The measure of the complement of an angle is 37+2x. The measure of a supplement of an angle is 141+x. What is the angle?
I solved for the two equations: c=37+2x and x+c=90. x = 17.666667
Then I solved for s=141+x and s+x=180. x = 19.5
The problem is that the angle is suppose to be the same.

Answer by stanbon(48516)

The measure of the complement of an angle is 37+2x. The measure of a supplement of an angle is 141+x. What is the angle?
--------------
37+2x=90-x
3x = 53
x = 53/3
--------------------
141+x = 180-x
2x = 39
x = 39/2
------------
I get the same results as you do.
================
Cheers,
Stan H.

I solved for the two equations: c=37+2x and x+c=90. x = 17.666667
Then I solved for s=141+x and s+x=180. x = 19.5